OpenStudy (anonymous):

[7.04] Choose the correct simplification of 5 to the 4th power over 5 to the 2nd power. Select one: a. 1 over 25 b. 10 c. 25 d. 1 over 10

OpenStudy (johnweldon1993):

\[\huge \frac{5^4}{5^2} = 5^{4 - 2} = 5^2 = \space ?\]

OpenStudy (anonymous):

25, so 'c'

OpenStudy (johnweldon1993):

Perfect :)

OpenStudy (anonymous):

Thank you love.

OpenStudy (johnweldon1993):

Anytime hun :)